show that any g€l(v, c) and u€v with g(u) not equal to 0 v=null g i{£u:£€c


1
Expert's answer
2021-10-07T14:08:28-0400

To show that V = null g ⊕ {"\lambda"u | "\lambda" ∈ C}, we need to show that V = null g + {"\lambda"u | "\lambda" ∈ C} and that null g ∩ {"\lambda"u | "\lambda" ∈ C} = {0}.

First we show that V = null g + {"\lambda"u | "\lambda" ∈ C}. Let "v\isin V". We need to find "n\isin null\ g" and

"w\isin" {"\lambda"u | "\lambda" ∈ C} for which v=n+w.  Suppose "g(v)=\lambda\isin C", and that "g(u)=\mu\isin C". We know that "\mu \neq 0" because u is not in null g. Thus, "\mu" has an inverse in C. Let "\alpha=\lambda \mu^{-1}\isin C". Note that this means "\alpha" times "\mu" gives back "\lambda" . So,

"g(\alpha u)=\alpha g(u)=\alpha\mu=\lambda"

Thus, "g(\alpha u)=g(v)". Let "n=v-\alpha u". Then "g(n)=g(v)-g(\alpha u)=0". Therefore, n ∈ null g. Set w = "\alpha"u. Then "w\isin" {"\lambda"u | "\lambda" ∈ C}, and v = n + w. So we can write v as the sum of elements of null g and {"\lambda"u | "\lambda" ∈ C}. Therefore, V = null g + {"\lambda"u | "\lambda" ∈ C}.

Next we show that null g ∩ {"\lambda"u | "\lambda" ∈ C} = {0}.  Suppose "v\isin" null g ∩ {"\lambda"u | "\lambda" ∈ C}. Then

"v\isin" {"\lambda"u | "\lambda" ∈ C}, so "v=\lambda u" for some "\lambda\isin C".

Since v ∈ null g, g(v) = 0. So g("\lambda"u) = 0. But g("\lambda"u) = "\lambda"g(u). Since "\lambda"g(u) = 0, either "\lambda" = 0 or g(u) = 0. But u is not in null g, so g(u) "\neq" 0. This means "\lambda" must equal 0. So v = "\lambda"u implies that v = 0. Therefore, null g ∩ {"\lambda"u | "\lambda" ∈ C} = {0}. Since V = null g + {"\lambda"u | "\lambda" ∈ C} and that

null g ∩ {"\lambda"u | "\lambda" ∈ C} = {0}, we have that V = null g ⊕ {"\lambda"u | "\lambda" ∈ C}.